OpenStudy (anonymous):

Add and simplify if possible:

OpenStudy (anonymous):

\[\frac{ 2 }{ 6x } + \frac{ 4 }{ 3x^2 }\]

OpenStudy (anonymous):

Please help :(

OpenStudy (ivettef365):

first find the common denominator

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x } + \frac{ 4 }{ 3x^2 } ?\]

OpenStudy (ivettef365):

you have 2 4 2(x) + 4(2) ---- + ----- = ----------------- 6x 3x^2 6x^2

OpenStudy (ivettef365):

oh, you can do it that way, then you have 1 4 1(x) + 4 ---- + ------ = ------------ 3x 3x^2 3x^2

OpenStudy (anonymous):

So my final answer is \[\frac{ x + 4 }{ 3x^2 }?\]

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

Awesome :) Thank you soooo much :)

OpenStudy (ivettef365):

yw :)